please help.
question: solve:5(x-4)^(2/3)=80
i cant really remember how to solve i thought maybe you substitute the (2/3) in for x but i still get lost.
im so sorry if i put this in the wrong category and if i did please try to help me the best way you can thank you

Answer by lwsshak3(11628)

solve:5(x-4)^(2/3)=80
divide both sides by 5
(x-4)^(2/3)=16
raise both sides to the (3/2) power
x-4=16^(3/2)
x=16^(3/2)+4
16^(3/2)=(√16)^3=4^3=64
x=64+4
x=68
..
check:
5(68-4)^(2/3)
=5(64)^(2/3)
=5*4^2
=5*16
=80